Grade 12th passAlgebra

Solve the following equations and check your results.
2x/3 + 1 = 7x/15 + 3

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

2x/3+1=7x/15+3
⇒2x/3–7x/15=3–1
⇒(10x–7x)/15=2
⇒3x=2×15
⇒3x=30
⇒x=30/3
⇒x=10
Putting the value of x in RHS and LHS we get,
L H S⇒2 x 10/3 + 1
= 23/3
RHS = 7 x 10/15 + 3
= 14/3 + 3
= 23/3
⇒LHS=RHS

Thanks